OpenStudy (anonymous):

What value of q makes the statement true? (q + 14) = 3 + (14)

Nnesha (nnesha):

3 + (14) = 3 + 14 so add them

OpenStudy (camerondoherty):

17 - 14 = q

Nnesha (nnesha):

 because there is only one to distribute bracket so 1 times 14 is 14